OpenStudy (anonymous):

@MathematicA barrel shaped like a cylinder is laid on its side and rolled up a ramp that is 111ft long. The barrel has a circular base that turns 37 times in being rolled up the ramp. What is the diameter of the circular base? Use the value 3.14 for Pi . Round your answer to the nearest tenth. Do not round any intermediate steps. @MathematicA barrel shaped like a cylinder is laid on its side and rolled up a ramp that is 111ft long. The barrel has a circular base that turns 37 times in being rolled up the ramp. What is the diameter of the circular base? Use the value 3.14 for Pi . Round your answer to the nearest tenth. Do not round any intermediate steps. @Mathematics

OpenStudy (anonymous):

\[\huge \pi D= 111/37 =3 \rightarrow D=\frac{3}{\pi}\]